Solve the system by elimination. If the system is inconsistent or has dependent equations, say so.
-7x + 5y = -20
-3x - 2y = 8

Answer by fractalier(6550)

From
-7x + 5y = -20
-3x - 2y = 8
let's multiply the top one by 2 and the bottom one by five...then add them...we get
-14x + 10y = -40
-15x - 10y = 40
now add them to get
-29x = 0 and
x = 0
then
y = -4
Solution is (0, -4)